OpenStudy (anonymous):

-4(x + 10) - 6 = -3(x - 2).

OpenStudy (goformit100):

Solve for X

OpenStudy (anonymous):

do you know how to do the distributive property for the brackets?

OpenStudy (goformit100):

yes

OpenStudy (anonymous):

not you lol

OpenStudy (anonymous):

Yes but I don't have anywhere to do my work nor a caculator

OpenStudy (anonymous):

just fins a scrap paper then tell me what you get... ORR you can use the draw function right under the block where you type your reply to do your work pretty messy tho

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

Yes I just want to make sure you follow along

OpenStudy (anonymous):

okay I will

OpenStudy (anonymous):

alright so let's start off by distributing the terms of this equation can you do that?

OpenStudy (anonymous):

uhhh show me how?

OpenStudy (anonymous):

So you have your equation -4(x + 10) - 6 = -3(x - 2) when you distribute... You multiply the outside term by the inside terms Of course each mathematical operator separates a term Soo distributing -4(x + 10) - 6 = -3(x - 2) You get... -4x - 40 - 6 = -3x + 6 Notice the sign changes...make sure you remember the common rule -*- = + +*+ = + -*+ = -

OpenStudy (anonymous):

you get the distribution part?

OpenStudy (anonymous):

okay so I need to solve -4x - 40 - 6 = -3x + 6 now?

OpenStudy (anonymous):

yes Why don't you add like terms before we start solving?

OpenStudy (anonymous):

you do know how to right?

OpenStudy (anonymous):

uh i think so. so i would add the x's and the whole numbers together then?

OpenStudy (anonymous):

yeup try that

OpenStudy (anonymous):

7x = 40?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

You have this -4x - 40 - 6 = -3x + 6 When you move the term to the other side of the equation you change the sign So now we move the terms to get the x terms on one side and the numbers on the other So... -4x + 3x = 6 + 6 + 40 -x = 52 x = -52

OpenStudy (anonymous):

ohh okay thank you so much

OpenStudy (anonymous):

yeah no problemm

OpenStudy (anonymous):

can you help me with the rest ?

OpenStudy (anonymous):

the rest? how many?

OpenStudy (anonymous):

ima start new questions haha

OpenStudy (anonymous):

oh they better not be the same :P